Let's consider the NORMAL (bell-shaped) CURVE. The EMPIRICAL RULE says roughly 68% of our data (in this case your grades) will be within one standard deviation of the mean. 95.44% will be within 2 SD's and 99.74% will be within 3 SD's. SO WHAT?
We are using points to determine grades in our class: 90-100 is an "A", 80-89 = "B", 70-79 = "C", 60-69 = "D" and <60 = "F".
The Empirical Rule tells us how many in our class of 30 students will be in each point range. TELL US HOW MANY this will be in each grade group, at least in theory

Answer by stanbon(57387)

Let's consider the NORMAL (bell-shaped) CURVE. The EMPIRICAL RULE says roughly 68% of our data (in this case your grades) will be within one standard deviation of the mean. 95.44% will be within 2 SD's and 99.74% will be within 3 SD's. SO WHAT?
We are using points to determine grades in our class: 90-100 is an "A", 80-89 = "B", 70-79 = "C", 60-69 = "D" and <60 = "F".
The Empirical Rule tells us how many in our class of 30 students will be in each point range. TELL US HOW MANY this will be in each grade group, at least in theory
--------
You have no mean grade and no standard deviation for the grades.
--------
If you assume the grades are between 60 and 100
you might say the mean grade is (100+60)/2 = 80
---
You are assuming the range of the grades is 60 to 100
so 6*sigma = 40. Then the standard deviation would
be 40/6 = 6.67
----
34% of 30 = 10.2
So 10 scores would be between 70 and 77
Also 10 scores would be between 70 and 63
---
You can figure out the rest.
Cheers,
Stan H.